OpenStudy (anonymous):

i need help please . i have to find the interior and the arc length of a pizza . which had a diameter of 16 and is able to be divided into 10 and 8 slices . please help

OpenStudy (allank):

@Cemantha Kindly elaborate. Do you want the arc length and centre angle of one slice of pizza for both cases? Case 1: when the pizza is divided into 10 slices. Case 2: when the pizza is divided into 8 slices

OpenStudy (anonymous):

yes please.

OpenStudy (allank):

I'll do the first case, then you should be able to do the second. Case 1: Diameter 16, divided into 10 slices. Find arc length and centre angle. Let's start with the centre angle. Keep in mind that the total angle at the centre is 360. Since the pizza is divided into equal slices, these 360 degrees are also divided equally, thus the centre angle = 360/10 = 36 degrees. Let's continue to the arc length. This can be directly related to the centre angle, in that, the length of the arc = (centre angle/360)*circumference Length of arc = (36/360)*22/7*16 = 5.0286 I hope that is clear enough and that you can do the second. :)

OpenStudy (anonymous):

oh ive done case 2 . can you tell me if its correct? Case 2: Diameter 16, divided into 8 slices. Find arc and centre angle. Lets start with the centre angle. Keep in mind that the totak angle at the centre is 360. Since the pizza is divide into equal slices, these 360 degrees are also divided equally, thus the centre andgle = 360/8 = 45 degrees. Lets continue to the arc length. This can be directly related to the centre angle, in that, the length of the arc = (centre angle/360)*circumference Length of arc = (45/360)*22/7*16 = 6.2857

OpenStudy (allank):

Wonderful. Very correct :)

OpenStudy (anonymous):

thank you. can yu help me find the area of each slice?

OpenStudy (allank):

You're welcome. Lemme give you a hint: area of slice = (centre angle/360)*area of whole pizza

OpenStudy (anonymous):

Area of slice = (36/360)*16 = 1.6

OpenStudy (anonymous):

for case 1 ^^^

OpenStudy (anonymous):

Area of slice = (45/360)*16 = 2 for case 2 ^^^

OpenStudy (allank):

remember that area of a pizza (circle) = pi*r^2

OpenStudy (allank):

Care to try again? you have the idea :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3.14*8^2 ->3.14*64 -> 200.96 is this correct?

OpenStudy (allank):

True. So that gives you the area of the pizza. continue...

OpenStudy (anonymous):

so it wud b (36/360)*200.96 = 20.096 and (45/360)*200.96 = 25.12

OpenStudy (allank):

Exactly =)

OpenStudy (anonymous):

thank you . soo much

OpenStudy (allank):

You're welcome.